if tan(A-B)=1,sec(A+B)=2/sq. root 3, then find the smallest positive valye of B?
 

Grade:11

1 Answers

Arun
6 years ago
/4)\pi/6-\pi+1/2(\pi/6)b=2n\pi/4+\pi+1/2(\pia=2n/6\pi+\pi/4 1/cos(a+b)=sec(a+b)=2/root 3cos(a+b)=root 3/2a+b=2n\pi+\pia-b=n/4)\pi/6-\pi/6)smallest value of b=1/2(\pi/4+\pifor n=0,smallest value of a=1/2(